NvidiaIntely (nvidiaintely):

A certain chemical reaction experiences a 7.50 x 10-7 kilogram loss in mass. If this mass were to be totally converted to energy, what would be the magnitude of that energy? (1 J = 1 kg m2/s2) 1.20 x 1023 J 2.5 x 109 J 4.00 x 101 J 6.75 x 1010 J

NvidiaIntely (nvidiaintely):

Thanks @Kevin for all your help!

OpenStudy (kevin):

E = m x c^2 E = 7.50*10^(-7) * (3*10^8)^2 E = 67500000000

NvidiaIntely (nvidiaintely):

Which is option D.

NvidiaIntely (nvidiaintely):

:)

OpenStudy (kevin):

sure :)